OpenStudy (anonymous):

WILL GIVE GOLD MEDAL Solve for n n - 1/3 = 5/3

OpenStudy (nurali):

n - 1/3 = 5/3 add both side by 1/3 n - 1/3 + 1/3 = 5/3+1/3 n = 5/3+1/3 taking LCM n= 5+1/3 n=6/3 n=2

OpenStudy (08surya):